# Properties and dynamics of generalized squeezed states

### v1

@@ I Introduction
Remarkable advances have been made in the past few decades in the field of quantum information technology Nielsen ; Fedorov ; Kimble ; Wehner ; Degen . Quantum computers with hundreds of qubits are already in commercial use. Quantum secure communication systems are being constructed around the world. Quantum sensing techniques have achieved precision that surpasses classical sensing limits. More progress on all these technologies is expected in the coming years.
@@
For many years, the study of nonlinear effects in light relied on combining strong input lasers with weak quadratic nonlinearities to generate weak but nonclassical output photon states. In recent years, especially with the development of circuit quantum electrodynamics in superconducting circuits, it has become possible to achieve much stronger nonlinearities and demonstrate various nonlinearity-based effects Yurke ; Movshovich ; Yamamoto ; Kudra . Implementations of nonlinear phenomena at to the single photon level are becoming increasingly common. Furthermore, it is now possible to access higher-order nonlinearities, e.g. cubic nonlinearities Chang ; Eriksson ; Saner2403 ; Saner2409 .
@@
One of the most extensively studied types of nonclassical states are squeezed states. Two-photon squeezed states, which can be obtained by using quadratic nonlinearities, have been investigated for decades. Although these states are nonclassical, they are still Gaussian, which limits their usefulness for quantum information processing. It is therefore important to investigate higher-order squeezed states, which are non-Gaussian and promise new applications in quantum computation and communication Ghose ; Hillery1999 ; DellAnno ; Zheng . In the past few years, there have been a few demonstrations of such higher-order squeezed states Chang ; Eriksson ; Saner2403 ; Saner2409 , as well as theoretical proposals for utilizing higher-order nonlinearities in physical setups Koshino ; Ayyash . While there have been some theoretical studies on this subject Fisher ; Hillery1984 ; Hong ; Braunstein1987 ; Braunstein1990 ; Hillery1990 ; Zelaya , as well as the related subject of three-photon down-conversion Banaszek ; Felbinger ; Olsen ; Bencheikh , a full picture of higher-order squeezing is lacking.
@@
In this work we aim to add insight to this area that is expected to play a role in future nonlinear quantum optics applications. We therefore perform a deeper analysis of the properties and dynamics of generalized squeezed states. We find some unexpected results, especially relating to the periodic dynamics, that differ qualitatively from what is known about the lower-order operations of displacement and two-photon squeezing. As the squeezing order is increased, the deviation of the quantum state from the initial state decreases and the quantum state returns to the initial state more rapidly. In other words, the maximum amount of squeezing that can be achieved decreases with increasing squeezing order.
@@
The remainder of this paper is organized as follows: in Sec. II, we provide some background information about previous work, and we describe our numerical simulation methodology. In Sec. III, we present the results of our numerical simulations, including sequences of Q and Wigner functions, as well as plots depicting the dependence of the photon number on the squeezing parameter. We discuss the interpretation of our numerical simulation results and their implications for physical realizations of generalized squeezing. In Sec. IV, we discuss the results further from a few additional points of view. In Sec. V, we end with some concluding remarks.
@@ II Background and methodology
It is now natural to consider the $n$th-order (or $n$-photon) generalization of displacement and two-photon squeezing:
@@
$\hat{H}_{n}=i\left[\left(\hat{a}^{\dagger}\right)^{n}-\hat{a}^{n}\right].$ (6)
@@
is a Hermitian operator. One can then use physics intuition to think about the squeezing operator by treating the squeezing parameter $r$ as the dimensionless time variable and treating $\hat{H}_{n}$ as the effective dimensionless Hamiltonian that generates the squeezing dynamics. From this point of view, it is straightforward to see that the operator $\hat{U}_{n}\left(r\right)$ is unitary and therefore cannot have any divergent matrix elements.
@@
In this work, we use numerical calculations to investigate the effect of the generalized squeezing operator. We perform a variety of checks to look for any indications that unphysical mathematical artefacts might be appearing in the numerical results. As we shall explain below, we do not see any such indications.
@@
Our calculations proceed as follows: we construct a large $N\times N$ matrix that represents a truncated version of the Hamiltonian in Eq. (6). Since all the matrix elements in $\hat{H}_{n}$ are equal to zero except between states that are different by exactly $n$ photons, we can divide the $N\times N$ matrix into $n$ smaller matrices of size $(N/n)\times(N/n)$, assuming that $N/n$ is an integer. Each one of these $n$ matrices corresponds to one subspace of the full Hilbert space. The $n$ orthogonal subspaces are $\left\{\left|0\right\rangle,\left|n\right\rangle,\left|2n\right\rangle,...\right\}$, $\left\{\left|1\right\rangle,\left|n+1\right\rangle,\left|2n+1\right\rangle,...\right\}$, …, $\left\{\left|n-1\right\rangle,\left|2n-1\right\rangle,\left|3n-1\right\rangle,...\right\}$. For example, for $n=3$, we have three subspaces that do not mix: $\left\{\left|0\right\rangle,\left|3\right\rangle,\left|6\right\rangle,...\right\}$, $\left\{\left|1\right\rangle,\left|4\right\rangle,\left|7\right\rangle,...\right\}$ and $\left\{\left|2\right\rangle,\left|5\right\rangle,\left|8\right\rangle,...\right\}$. This separation of the full $N\times N$ matrix into $n$ smaller ones simplifies further manipulation of the matrices. Once we have constructed one of the relevant matrices that represents $\hat{H}_{n}$, we evaluate the corresponding squeezing operator for a small value of the squeezing parameter $\delta r$ by taking the exponential of the truncated Hamiltonian matrix multiplied by $-i\delta r$. Specifically, we always use $\delta r=0.01$. Once we have $\hat{U}_{n}(\delta r)=\exp\left\{-i\delta r\hat{H}_{n}\right\}$, we evaluate the squeezing operator for larger values of $r$ by multiplying $\hat{U}_{n}(\delta r)$ by itself $r/\delta r$ times. In other words, $\hat{U}_{n}(r)=\left[\hat{U}_{n}(\delta r)\right]^{r/\delta r}$. We performed a few random checks by evaluating $\hat{U}_{n}(r)$ in one step as $\exp\left\{-ir\hat{H}_{n}\right\}$ and inspecting the amplitudes in the state $\hat{U}_{n}(r)\left|0\right\rangle$. We found that the results differed from those obtained when using the formula $\hat{U}_{n}(r)=\left[\hat{U}_{n}(\delta r)\right]^{r/\delta r}$ by only $\sim 10^{-10}$, which is an indication that we are not encountering any serious issues with the numerical evaluation of $\exp\left\{-ir\hat{H}_{n}\right\}$. Since some previous studies that performed numerical simulations of generalized squeezing had the issue of an unconserved total probability, we point out that in our simulations the quantum state always remained normalized, as it should. Using the squeezing operator for a given value of $r$, or for a sequence of $r$ values, we can analyze various properties of the squeezed states as well as the evolution of these properties as functions of the squeezing parameter $r$.
@@
where we have moved the index $n$ to the superscript to avoid having a subscript with three indices. An important factor in giving rise to divergences is the fact that, for large $n$, the matrix elements grow very fast for large $k$. As a result, we must look for potential issues related to the truncation of the Hamiltonian matrix. One check that we perform for this purpose is that we investigate the effect of the specific value of the matrix size $N$ used in our calculations. In other words, we vary the size of the matrix $N$ and compare the results of a few different calculations with different values of $N$. As a further check, we construct soft-cutoff matrices in which we take the original (hard-cutoff) $N\times N$ Hamiltonian matrix and multiply the matrix elements in the bottom-right corner of the matrix by a suppression factor such that the matrix element decrease and approach zero in the bottom-right corner of the matrix. Specifically, when $(N-k)/N<0.2$, we multiply the matrix element $H_{k,k+n}^{(n)}$ by the factor $\sin[5\pi(N-k)/(2N)]$, which decreases from 1 at $k=0.8N$ to zero at $k=N$. The matrix element dependence on $k$ and cutoff are illustrated in Fig. 1.
@@
Our approach of comparing different matrices that all approximate $\hat{H}_{n}$ can also be interpreted differently from just being a check on whether our numerical approach is sound or not. From a practical point of view, one can say that, while the behaviour of an infinite perfect system is of pure theoretical interest, it is important to identify results that are robust to variations in the Hamiltonian away from the perfect form given in Eq. (6). In a practical situation, e.g. in a superconducting circuit containing nonlinear elements whose purpose is to induce the generalized squeezing effect, we do not expect that the physical system will be accurately described by the Hamiltonian in Eq. (6) beyond tens of photons. A system that contains a cubic nonlinearity will almost invariably contain quartic and higher-order nonlinearities whose effects become increasingly nonnegligible with increasing photon number. These effects can naturally impose a cutoff on the number of photons up to which Eq. (6) is a good approximation.
@@ III Results
Figure 2: Q and Wigner functions for the trisqueezed vacuum state, i.e. $\hat{U}_{3}(r)\left|0\right\rangle$, for $r=0.1$ (top-left), 0.5 (top-right) and 0.7 (middle-left), $r=0.86$ (middle-right), 1 (bottom-left) and 1.5 (bottom-right). In all Q and Wigner function plots in this manuscript, the $x$ and $y$ axes correspond, respectively, to the position ($x$) and momentum ($p$) variables of phase space. The range of the axes is $[-5,5]$.
@@
Figure 4: Occupation probabilities of the lowest 11 relevant states ($\left|0\right\rangle$, $\left|3\right\rangle$, $\left|6\right\rangle$, … $\left|30\right\rangle$) for the trisqueezed vacuum state with $r=1$, i.e. $\hat{U}_{3}(r=1)\left|0\right\rangle$, as functions of photon number $m$. Each value of $m$ has eight different data points. The ones shifted slightly to the left correspond to a hard cutoff, while the ones shifted slightly to the right correspond to a soft cutoff. The red dots, green squares, blue + signs and magenta $\times$ symbols correspond, respectively to $N=300$, 3000, $3\times 10^{4}$ and $6\times 10^{4}$. The results are almost independent of the truncation details except for two points, namely the states $\left|0\right\rangle$ and $\left|3\right\rangle$. We performed additional simulations with $N=1.2\times 10^{4}$, $N=2.1\times 10^{4}$ and $N=4.5\times 10^{4}$, and the results (not shown) were very close to those with $N=3\times 10^{4}$ and $N=6\times 10^{4}$.
Figure 5: Q and Wigner functions for the four-photon squeezed vacuum state, i.e. $\hat{U}_{4}(r)\left|0\right\rangle$, for $r=0.1$ (top-left), 0.2 (top-right) and 0.3 (middle-left), 0.37 (middle-right), 0.5 (bottom-left) and 0.6 (bottom-right).
@@
First, we performed numerical simulations following the method described in Sec. II for the exactly solvable cases of displacement ($n=1$) and two-photon squeezing ($n=2$). Our results consistently agreed with the known solutions for these two cases. We will show only a few results of these cases below, mainly for the purpose of comparing with the case $n\geq 3$.
@@
Next we performed simulations for the case of trisqueezing, i.e. $n=3$. In Fig. 2, we plot the Q and Wigner functions for the trisqueezed vacuum state, i.e. $\hat{U}_{3}(r)\left|0\right\rangle$, for six values of the squeezing parameter $r$. The Q and Wigner functions serve complementary purposes. The Q function shows most clearly the coherent states that participate in a quantum superposition of such states. Note that while the appearance of the Q function suggests that it contains only probability information, i.e. with no phase information about superpositions of coherent states, it does in fact contain (in its detailed and often not-easily-discernible features) the full information about the quantum state. On the other hand, the Wigner function tends to be nonzero in a larger area of phase space, but it conveys the presence of non-Gaussian features prominently in the form of negative values of the function.
@@
For small values of the squeezing parameter $r$ (Fig. 2 top-right), the Q and Wigner functions have a central peak at the origin of phase space with three small arms extending along three directions that make 120 degree angles with each other, as is intuitively expected and as was shown in previous studies Braunstein1987 ; Braunstein1990 . The Wigner function takes negative values in curved regions along the three sides of the main peak with its extensions. Another feature that we can see in the Q function are three broad and faint peaks along the directions between the three main arms. As we increase the squeezing parameter further (Fig. 2 middle-left), the central peak is greatly suppressed, especially as seen in the Q function, and we instead have a pattern that looks like that of a superposition of three coherent states, somewhat similar to the one synthesized in Ref. Hofheinz . As we increase $r$ further, the three peaks do not continue to move farther from the origin. Instead, their height decreases, and the three peaks that were faint for small $r$ increase in height. At $r=0.86$ (Fig. 2 middle-right), the six peaks seem to have almost exactly the same height and shape. As we increase $r$ further (Fig. 2 bottom-left), the three peaks that were faint become the dominant ones, while the three peaks that were dominant become faint. As we increase $r$ further (Fig. 2 bottom-right), the three peaks approach the origin, merge and return almost completely to the initial state, i.e. the vacuum state.
@@
The periodic behaviour can also be seen clearly if we plot the occupation probabilities of the different photon number states as functions of $r$, as shown in Fig. 3. While the oscillations are not exactly symmetric in $r$, the occupation probability of the vacuum state goes back very close to its initial value at some point. Specifically, at $r=1.75$ the probability is 0.991. On the other hand, the occupation probability of the vacuum state reaches its lowest value (below $5\times 10^{-4})$ at 0.9, which is slightly shifted from the point ($r=0.86$) where the Q and Wigner functions are closest to having six-fold symmetry. Note that in Fig. 3 the highest probability Fock state alternates between being $\left|0\right\rangle$ or $\left|3\right\rangle$. For comparison, the probability distribution for ($n=1$) keeps moving to higher photon numbers indefinitely, while for $n=2$ the probability is a monotonically decreasing function (considering only even photon numbers) and the highest probability Fock state is always $\left|0\right\rangle$.
@@
Next we turn to the case of four-photon squeezing, i.e. $n=4$. The Q and Wigner functions are shown in Fig. 5. Similarly to what we found in the case of trisqueezing, the Q function develops four small arms along the $x$ and $p$ axes in phase space. For larger values of $r$, the Q function resembles a superposition of four coherent states centered at symmetrically located points on the $x$ and $p$ axes. These peaks then gradually fade and are replaced by four peaks located symmetrically along the diagonal lines between the $x$ and $p$ axes. The Q and Wigner functions then evolve back to the vacuum state. From the plot of the occupation probabilities as functions of $r$, shown in Fig. 6, we find that the maximum vacuum state probability after the first oscillation is $0.9989$, and it occurs at $r=0.75$. It is interesting that now the occupation probabilities with $N=400$ and $N=4\times 10^{4}$ are almost indistinguishable.
@@
The average photon number $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ for the state $\hat{U}_{n}(r)\left|0\right\rangle$, plotted as a function of the squeezing parameter $r$ in Fig. 7, exhibits a few interesting features. For comparison, we include in Fig. 7 simulation results for $n=1$ (displacement) and $n=2$ (two-photon squeezing). In these two cases, the average photon number grows indefinitely. For trisqueezing ($n=3$), the average photon number exhibits irregular oscillations. As we progressively increase $n$ to $n=4$ and $n=5$, the oscillations become less irregular. More specifically, the oscillations become increasingly close to simple sinusoidal oscillations. Furthermore, the maximum value of $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ decreases with increasing $n$. The oscillation period also decreases with increasing $n$. A prominent feature in Fig. 7 relates to the height of the oscillations. The maximum height for $n=3$ reaches hundreds of photons, which is much higher than what we might expect from Figs. 2-4. It is the small probability amplitude that goes close to the maximum allowed photon number, i.e. $N\sim 10^{4}$, that results in an average photon number on the order of hundreds. To investigate this point further, in Fig. 8 we plot the maximum value of $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ reached in the first oscillation as a function of matrix size $N$. For $n=3$, by fitting the data in Fig. 8(a) to a straight line, we find that to a good approximation $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle_{\rm max}\propto N^{0.562}$ for the case of a hard cutoff and $N^{0.610}$ for the case of a soft cutoff. These functions diverge as $N\to\infty$. However, it should be noted that most of the probability remains at small values of $m$, and it is the small probability that goes to large photon number states that raises the average photon number, possibly up to infinity. Another remarkable feature in this context is the fact, although the height of the oscillations in Fig. 7(b) varies greatly for the different curves, the overall pattern seems to be almost independent of the truncation details. This point is illustrated most clearly in the fact that, by coincidence, two of the curves (which have different truncation conditions) seem to overlap perfectly. Furthermore, the point at which $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ goes back close to zero, i.e. the period of the oscillations, seems to be the same for all the curves in the figure.

### v2

@@ I Introduction
Remarkable advances have been made in the past few decades in the field of quantum information technology Nielsen ; Fedorov ; Kimble ; Wehner ; Illiano ; Degen . Quantum computers with hundreds of qubits are already in commercial use. Quantum secure communication systems are being constructed around the world. Quantum sensing techniques have achieved precision that surpasses classical sensing limits. More progress on all these technologies is expected in the coming years.
@@
For many years, the study of nonlinear effects in light relied on combining strong input lasers with weak quadratic nonlinearities to generate weak but nonclassical output photon states. In recent years, especially with the development of circuit quantum electrodynamics in superconducting circuits, it has become possible to achieve much stronger nonlinearities and demonstrate various nonlinearity-based effects Yurke ; Movshovich ; Yamamoto ; Kudra . Implementations of nonlinear phenomena at to the single photon level are becoming increasingly common. Furthermore, it is now possible to access higher-order nonlinearities, e.g. cubic nonlinearities Chang ; Eriksson ; Bazavan ; Saner .
@@
One of the most extensively studied types of nonclassical states are squeezed states. Two-photon squeezed states, which can be obtained by using quadratic nonlinearities, have been investigated for decades. Although these states are nonclassical, they are still Gaussian, which limits their usefulness for quantum information processing. It is therefore important to investigate higher-order squeezed states, which are non-Gaussian and promise new applications in quantum computation and communication Ghose ; Hillery1999 ; DellAnno ; Zheng . For example, higher-order squeezing could be used to generate multi-photon entangled states DellAnno . Furthermore, combining a non-Gaussian operation with displacement and squeezing can be a way to achieve universal control of photonic states Eriksson . In the past few years, there have been a few demonstrations of higher-order squeezed states Chang ; Eriksson ; Bazavan ; Saner , as well as theoretical proposals for utilizing higher-order nonlinearities in physical setups Koshino ; Ayyash . While there have been some theoretical studies on this subject Fisher ; Hillery1984 ; Hong ; Braunstein1987 ; Braunstein1990 ; Hillery1990 ; Zelaya ; Braak , as well as the related subject of three-photon down-conversion Banaszek ; Felbinger ; Olsen ; Bencheikh , a full picture of higher-order squeezing is lacking.
@@
In this work we aim to add insight to this area that is expected to play a role in future nonlinear quantum optics applications. We therefore perform a systematic analysis of the properties and dynamics of generalized squeezed states. We find some unexpected results, especially relating to periodic dynamics, that differ qualitatively from what is known about the lower-order operations of displacement and two-photon squeezing. As the squeezing order is increased, the deviation of the quantum state from the initial state decreases and the quantum state returns to the initial state more rapidly. In other words, the maximum amount of squeezing that can be achieved decreases with increasing squeezing order.
@@
The remainder of this paper is organized as follows: in Sec. II, we provide some background information about previous work, and we describe our numerical simulation methodology. In Sec. III, we present the results of our numerical simulations, including sequences of Husimi and Wigner functions, as well as plots depicting the dependence of the photon number on the squeezing parameter. We discuss the interpretation of our numerical simulation results and their implications for physical realizations of generalized squeezing. In Sec. IV, we discuss the results from a few additional points of view. In Sec. V, we end with some concluding remarks.
@@ II Background and methodology
It is then natural to consider the $n$th-order (or $n$-photon) generalization of displacement and two-photon squeezing:
@@
$\hat{H}_{n}=i\left[\left(\hat{a}^{\dagger}\right)^{n}-\hat{a}^{n}\right]$ (6)
@@
is a Hermitian operator. One can then use physics intuition to think about the squeezing operator by treating the squeezing parameter $r$ as the dimensionless time variable and treating $\hat{H}_{n}$ as the effective dimensionless Hamiltonian that generates the squeezing dynamics. From this point of view, it is straightforward to conclude that the operator $\hat{U}_{n}\left(r\right)$ is unitary and therefore cannot have any divergent matrix elements.
@@
In this work, we use numerical calculations to investigate the effect of the generalized squeezing operator on the vacuum state. We perform a variety of checks to look for any indications that unphysical mathematical artefacts might be appearing in the numerical results. As we shall explain below, we do not see any such indications.
@@
Our calculations proceed as follows: we construct a large $N\times N$ matrix that represents a truncated version of the Hamiltonian in Eq. (6). Since all the matrix elements in $\hat{H}_{n}$ are equal to zero except between states that are different by exactly $n$ photons, we can divide the $N\times N$ matrix into $n$ smaller matrices of size $(N/n)\times(N/n)$, assuming that $N/n$ is an integer. Each one of these $n$ matrices corresponds to one subspace of the full Hilbert space. The $n$ orthogonal subspaces are $\left\{\left|0\right\rangle,\left|n\right\rangle,\left|2n\right\rangle,...\right\}$, $\left\{\left|1\right\rangle,\left|n+1\right\rangle,\left|2n+1\right\rangle,...\right\}$, …, $\left\{\left|n-1\right\rangle,\left|2n-1\right\rangle,\left|3n-1\right\rangle,...\right\}$. For example, for $n=3$, we have three subspaces that do not mix: $\left\{\left|0\right\rangle,\left|3\right\rangle,\left|6\right\rangle,...\right\}$, $\left\{\left|1\right\rangle,\left|4\right\rangle,\left|7\right\rangle,...\right\}$ and $\left\{\left|2\right\rangle,\left|5\right\rangle,\left|8\right\rangle,...\right\}$. This separation of the full $N\times N$ matrix into $n$ smaller ones simplifies further manipulation of the matrices. Once we have constructed one of the relevant matrices that represent $\hat{H}_{n}$, we evaluate the corresponding squeezing operator for a small value of the squeezing parameter $\delta r$ by taking the exponential of the truncated Hamiltonian matrix multiplied by $-i\delta r$, i.e. we evaluate $\hat{U}_{n}(\delta r)=\exp\left\{-i\delta r\hat{H}_{n}\right\}$. We always use $\delta r=0.01$. We evaluate the squeezing operator for larger values of $r$ by multiplying $\hat{U}_{n}(\delta r)$ by itself $r/\delta r$ times. In other words, $\hat{U}_{n}(r)=\left[\hat{U}_{n}(\delta r)\right]^{r/\delta r}$. We performed a few random checks by evaluating $\hat{U}_{n}(r)$ in one step as $\exp\left\{-ir\hat{H}_{n}\right\}$ and inspecting the amplitudes in the state $\hat{U}_{n}(r)\left|0\right\rangle$. We found that the results differed from those obtained when using the formula $\hat{U}_{n}(r)=\left[\hat{U}_{n}(\delta r)\right]^{r/\delta r}$ by only $\sim 10^{-10}$, which is an indication that we are not encountering any serious issues with the numerical evaluation of $\exp\left\{-ir\hat{H}_{n}\right\}$. Since some previous studies that performed numerical simulations of generalized squeezing had the issue of an unconserved total probability, we point out that in our simulations the quantum state always remained normalized, as it should. We vary the squeezing parameter $r$, evaluate $\hat{U}_{n}(r)\left|0\right\rangle$ and analyze the state as a functions of $r$.
@@
where we have moved the index $n$ to the superscript to avoid having a subscript with three indices. An important factor in giving rise to divergences is the fact that, for large $n$, the matrix elements grow very fast for large $k$. As a result, we must be careful and try to analyze how our results depend on the truncated matrix size, so that we can make predictions about the results for an infinite, untruncated matrix. For this purpose, we perform several calculations with the same system parameters but with different matrix sizes, and we investigate how our results change as we vary the matrix size $N$. As a further check on the effect of matrix truncation on our results, we construct soft-cutoff matrices in which we take the original (hard-cutoff) $N\times N$ Hamiltonian matrix and multiply the matrix elements in the bottom-right corner of the matrix by a suppression factor, such that the matrix elements in the bottom-right corner of the matrix decrease and approach zero smoothly. The specific suppression factor that we use is equal to 1 for $k/N<0.8$, such that all matrix elements with $k<0.8N$ remain unchanged. However, when $k/N>0.8$, we multiply the matrix element $H_{k,k+n}^{(n)}$ by the factor $\sin[5\pi(N-k)/(2N)]$, which decreases from 1 at $k=0.8N$ to zero at $k=N$. The matrix element dependence on $k$ for the cases of hard and soft cutoff are illustrated in Fig. 1. In the case of the hard cutoff, the matrix elements keep increasing all the way up to the highest value of $k$ kept in the truncated matrix. This situation is analogous to having a hard wall when studying the dynamics of a moving particle. In the case of the soft cutoff, the matrix elements decrease smoothly as we approach the high-$k$ edge of the truncated matrix. We note that the details of our soft-cutoff suppression factor are arbitrary. We expect that our conclusions will remain the same for any generally similar soft-cutoff truncation. In fact, we performed a few test simulations in which the soft cutoff was introduced by adding gradually increasing diagonal matrix elements in the bottom-right corner of $\hat{H}_{n}$. The results of these simulations were qualitatively similar to the results with the other cutoff conditions that we used in this study.
@@
Our approach of comparing different matrices that all approximate $\hat{H}_{n}$ can also be interpreted differently from just being a check on whether our numerical approach is sound or not. From a practical point of view, one can say that, while the behaviour of an infinite perfect system is of pure theoretical interest, it is important to identify results that are robust to variations in the Hamiltonian away from the perfect form given in Eq. (6). In a practical situation, e.g. in a superconducting circuit containing nonlinear elements whose purpose is to induce the generalized squeezing effect, we do not expect that the physical system will be accurately described by the Hamiltonian in Eq. (6) beyond tens of photons. A system that contains a cubic nonlinearity will almost invariably contain quartic and higher-order nonlinearities whose effects become increasingly nonnegligible with increasing photon number. These effects can naturally impose a cutoff on the number of photons up to which Eq. (6) is a good approximation. From this point of view, it might make more sense to use the finite-$N$ simulation results as a reference point to test and judge the soundness and/or relevance of infinite-$N$ treatments, rather than the other way around.
It is also worth mentioning here that the Hamiltonian in Eq. (6) is unbounded from below, i.e. it predicts infinite negative energies, and can therefore be considered unphysical. One must note, however, that this Hamiltonian is typically obtained as an effective Hamiltonian in a rotating frame. If the initial state has a small number of photons, the dynamics generated by the Hamiltonian will not take the system to the infinite-photon-number region, where the unphysical effects are encountered. It should be noted that the issue of boundedness applies also for the cases $n=1$ and $n=2$, where it is known that this Hamiltonian, or truncated versions of it, give accurate descriptions of the dynamics.
@@ III Results
Figure 2: Q and Wigner functions for the trisqueezed vacuum state, i.e. $\hat{U}_{3}(r)\left|0\right\rangle$, for $r=0.1$, 0.5, 0.7, 0.86, 1 and 1.5. In all Q and Wigner function plots in this manuscript, the $x$ and $y$ axes correspond, respectively, to the position ($x$) and momentum ($p$) variables of phase space. The range of the axes is $[-5,5]$.
@@
Figure 4: Occupation probabilities of the lowest 11 relevant states ($\left|0\right\rangle$, $\left|3\right\rangle$, $\left|6\right\rangle$, … $\left|30\right\rangle$) for the trisqueezed vacuum state with $r=1$, i.e. $\hat{U}_{3}(r=1)\left|0\right\rangle$, as functions of photon number $k$. Each value of $k$ has eight different data points. The ones shifted slightly to the left correspond to a hard cutoff, while the ones shifted slightly to the right correspond to a soft cutoff. The red dots, green squares, blue + signs and magenta $\times$ symbols correspond, respectively to $N=300$, 3000, $3\times 10^{4}$ and $6\times 10^{4}$. The results are almost independent of the truncation details except for two points, namely the states $\left|0\right\rangle$ and $\left|3\right\rangle$. We performed additional simulations with $N=1.2\times 10^{4}$, $N=2.1\times 10^{4}$ and $N=4.5\times 10^{4}$, and the results (not shown) were very close to those with $N=3\times 10^{4}$ and $N=6\times 10^{4}$.
Figure 5: Q and Wigner functions for the four-photon squeezed vacuum state, i.e. $\hat{U}_{4}(r)\left|0\right\rangle$, for $r=0.1$, 0.2, 0.3, 0.37, 0.5 and 0.6.
@@
First, we performed numerical simulations following the method described in Sec. II for the exactly solvable cases of displacement ($n=1$) and two-photon squeezing ($n=2$). Our results consistently agreed with the known solutions for these two cases. For example, when we set $r=2$ and $N=n\times 10^{4}$, the simulations gave average photon numbers that deviate from the exact values by $7\times 10^{-14}$ and $6\times 10^{-4}$ for the cases $n=1$ and $n=2$, respectively. This excellent agreement between the simulations and exact results serves as an additional check on the validity of the simulations with the typical parameters that we use. We will show only a few further results for the cases $n=1$ and $n=2$ below, mainly for the purpose of comparing with the case $n\geq 3$.
@@
Next we performed simulations for the case of trisqueezing, i.e. $n=3$. In Fig. 2, we plot the Q and Wigner functions for the trisqueezed vacuum state, i.e. $\hat{U}_{3}(r)\left|0\right\rangle$, for six representative values of the squeezing parameter $r$. The Q and Wigner functions serve complementary purposes. The Q function shows most clearly the coherent states that participate in a quantum superposition of such states. Note that while the appearance of the Q function suggests that it contains only probability information, i.e. with no phase information about superpositions of coherent states, it does in fact contain (in its detailed and often not-easily-discernible features) the full information about the quantum state. On the other hand, the Wigner function tends to be nonzero in a larger area of phase space, but it conveys the presence of non-Gaussian features prominently in the form of negative values of the function.
@@
For small values of the squeezing parameter $r$ ($r=0.1$ and 0.5), the Q and Wigner functions have a central peak at the origin of phase space with three small arms extending along three directions that make 120 degree angles with each other. This behaviour is what one would intuitively expect by considering the well-known two-photon squeezing behaviour along with the fact that $n$-th order squeezed vacuum states must have $n$-fold rotation symmetry, i.e. they must remain unchanged when rotated by an angle of $2\pi/n$. In other words, $\hat{R}^{\dagger}(2\pi/n)\hat{H}_{n}\hat{R}(2\pi/2n)=\hat{H}_{n}$, where $\hat{R}(x)=\exp\left\{ix\hat{a}^{\dagger}\hat{a}\right\}$, as explained in Ref. Braunstein1987 . In fact, similar Q functions appear in Refs. Braunstein1987 ; Braunstein1990 . The Wigner function takes negative values in the regions adjacent to the main peak between the directions of the three arms. Another feature that we can see in the Q function at $r=0.5$ are three broad and faint peaks along the directions between the three main arms. As we increase the squeezing parameter further ($r=0.7$), the central peak is greatly suppressed, especially as seen in the Q function, and we instead have a pattern that looks like that of a superposition of three coherent states, somewhat similar to the one synthesized in Ref. Hofheinz . As we increase $r$ further, the three peaks do not continue to move farther from the origin. Instead, their height decreases, and the three peaks that were faint for small $r$ increase in height. At $r=0.86$, the six peaks seem to have almost exactly the same height and shape. As we increase $r$ further ($r=1$), the three peaks that were faint become the dominant ones, while the three peaks that were dominant become faint. As we increase $r$ further ($r=1.5$), the three peaks approach the origin and merge into a single peak. At $r=1.75$ (not shown in Fig. 2) the state returns almost completely to the initial state, i.e. the vacuum state.
@@
The periodic behaviour can also be seen clearly if we plot the occupation probabilities of the different photon number states as functions of $r$, as shown in Fig. 3. While the oscillations are not exactly symmetric in $r$, the occupation probability of the vacuum state goes back very close to its initial value at some point. Specifically, at $r=1.75$ the probability is 0.991. On the other hand, the occupation probability of the vacuum state reaches its lowest value (below $5\times 10^{-4})$ at 0.9, which is slightly shifted from the point ($r=0.86$) where the Q and Wigner functions are closest to having six-fold symmetry. Note that in Fig. 3 the highest probability Fock state alternates between being $\left|0\right\rangle$ and $\left|3\right\rangle$. For comparison, the probability distribution for ($n=1$) keeps moving to higher photon numbers indefinitely, while for $n=2$ the probability is a monotonically decreasing function (considering only even photon numbers) and the highest probability Fock state is always $\left|0\right\rangle$.
@@
Next we turn to the case of four-photon squeezing, i.e. $n=4$. The Q and Wigner functions are shown in Fig. 5. Similarly to what we found in the case of trisqueezing, the Q function develops four small arms along the $x$ and $p$ axes in phase space. For larger values of $r$, the Q function resembles a superposition of four coherent states centered at symmetrically located points on the $x$ and $p$ axes. These peaks then gradually fade and are replaced by four peaks located symmetrically along the diagonal lines between the $x$ and $p$ axes. The Q and Wigner functions then evolve back to the vacuum state. From the plot of the occupation probabilities as functions of $r$, shown in Fig. 6, we find that the maximum vacuum state probability at the end of the first oscillation is $0.9989$, and it occurs at $r=0.75$. It is interesting that now the occupation probabilities with $N=400$ and $N=4\times 10^{4}$ (with different types of cutoff) are almost indistinguishable.
@@
The average photon number $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ for the state $\hat{U}_{n}(r)\left|0\right\rangle$, plotted as a function of the squeezing parameter $r$ in Fig. 7, exhibits a few interesting features. For comparison, we include in Fig. 7 simulation results for $n=1$ (displacement) and $n=2$ (two-photon squeezing). In these two cases, the average photon number grows indefinitely. For trisqueezing ($n=3$; shown in Fig. 7b), the average photon number exhibits irregular oscillations. A remarkable feature in the case $n=3$ is the fact that, although the height of the oscillations in Fig. 7(b) varies greatly for the different curves (which correspond to different truncation conditions), the overall pattern seems to be almost independent of the truncation details. This point is illustrated most clearly in the fact that, by coincidence, two of the curves (which have different values of $N$ and differ in whether they have a hard or soft cutoff) seem to overlap perfectly. Furthermore, the point at which $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ goes back close to zero, i.e. the period of the oscillations, seems to be the same for all the curves in the figure. We will come back to this point shortly. As we progressively increase $n$ to $n=4$ and $n=5$, the oscillations become less irregular. More specifically, the oscillations become increasingly close to simple sinusoidal oscillations. Furthermore, the maximum value of $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ decreases with increasing $n$. The oscillation period also decreases with increasing $n$. A prominent feature in Fig. 7 relates to the height of the oscillations. The maximum height for $n=3$ reaches hundreds of photons. At first sight, this value seems to contradict the results plotted in Figs. 2-4, which indicate that the probability distribution remains largely confined to small photon numbers. However, the long tail of the probability distribution that extends to the maximum photon number in the simulation, i.e. $N\sim 10^{4}$, is responsible for raising the average photon number to the order of hundreds.